OpenStudy (anonymous):

How to find the area under this curve: x = 1/2, x = sqrt(y) and y = -x+2? I can't visualize the graph.

OpenStudy (anonymous):

Have you learn't calculus

OpenStudy (anonymous):

Yes, it's a exercise of Integral Calculus.

OpenStudy (anonymous):

So you don't have to visualize anythng

OpenStudy (anonymous):

What is the upper limit and lower limit ?

hartnn (hartnn):

x = sqrt(y) parabola (upper part since it is positive sqrt y) and y = -x+2 line

OpenStudy (anonymous):

yes, i got it. My problem is I can't understand that x = sqrt(y). How do I write the integral?

hartnn (hartnn):

i mean 1st quadrant , since x is positive

hartnn (hartnn):